a particle starting fom mean position executes shm with a period of 8sec.The minimum time in which its potential energy becomes half of the total energy is

Grade:12th pass

1 Answers

Eshan
5 years ago
Time period of SHM=\dfrac{2\pi}{\omega}=8 seconds\implies \omega=\dfrac{\pi}{4}Hz

Potential energy of particle is half of the total energy.

\implies \dfrac{1}{2}kx^2=\dfrac{1}{2}\times \dfrac{1}{2}kA^2
Hence position at which the potential energy is half of total energy isx=\dfrac{A}{\sqrt{2}}

x=Asin(\omega t )=\dfrac{A}{\sqrt{2}}
\implies \omega t=\dfrac{\pi}{4} \implies t=1s